Project

General

Profile

2017-11-24-MIPO

Phone details - http://mice.iit.edu/phonebridge.html

RAL R1 2nd Floor East MICE Meeting room 15.30pm UK time, 09.30 Chicago

Notes

Present: KL, CR, JB, CW

  • Actions:
    • None recorded.
  • Magnet status - JB
    • Ramped to SSD-on-2T setting, with no issues. TDKs set so that no changes need to be made for ramp on Sunday. J.Cobb will be in attendance. Expect to have new setting on Sunday by 2pm.
    • On Monday morning D.Rajaram will ramp magnets down so that the PRY can be disassembled first thing in the am.
    • Short-term pressure rises observed on SSD, but always recover. Pressure excursions observed in previous operation. JB believes this to be mechanical issues inside the cold-head leading to fluctuations in temperature that lead to burps of gas.
  • Tracker
    • No present issue.
    • Need to decide whether to swap back the water to tracker cryostats. Pressure in the water circuit is now higher so cryostats are running colder. Focus-coil compressors will be moved back onto the proper circuit on Monday or Tuesday next week.
  • Controls - AK
    • Ajit on night shift from this evening. He has been responding to issues, but no particular issue to raise here.
    • JB reports several UPS alarms (warnings). Perhaps A.Kurup could check the origin of the UPS warnings. KL Send not to A.Kurup to alert him.
  • ISIS 2017/03 - 14th Nov to 20th Dec - CW/KL
    • Discussion of run plan and details of programme. Agreed to do flip mode SSD on/off first. Then change to solenoid mode.
    • CR: Will define settings for next block data to MIPO/MEMO for Monday 27Nov17.
  • US equipment: CW
    • Status remains as before. Need to respond to DL email re equipment at DL.
    • S.Hayward is preparing email to MICE management re definition of programme for moth-balling and use of MICE Muon Beam as test beam.
  • S/w&C report: DR
    • D.Rajaram on shift so no report.
  • Paper status: CR
    • Emittance analysis (V.Blackmore) making progress, now moving to referees meeting in mid January.
    • Scattering paper has not moved too much since J.Nugent is on shift.
    • Emittance change paper, CR is working on details of beam transport and energy lots.
    • Detailed survey of equipment inside the spectrometer solenoids needs to be planned. Action CW: prepare plan for post-run survey.
    • LH2 h/w paper initiated. Hardware paper to be initiated soon.
  • AOB

Documents

Phone Details

UK: 08082380274
US: 631.621.5253
Access code: 1779339517

2017-03-RunPlan-v10.xlsx (22 KB) Long, Kenneth, 24 November 2017 16:05